Suppose X a population has a distribution that is skewed to the right and µ =400 and σ =24. If a random sample of size 144 is drawn from the population:

   a) What is the name of theorem that allows us to solve this problem? ___________

   b) What is the reason that we use that theorem in this case? ___________

   c) Compute P(399 <  < 403) _________

In part c,   is the sample mean.

PLEASE SHOW ALL YOUR WORK
Answer #1

Solution :

Given that,

mean = \mu = 400

standard deviation = \sigma = 24

n = 144

a) Central Limit Theorem

b) because sample size is greater than or equal 30

\mu\bar x = \mu = 400

\sigma\bar x = \sigma / \sqrt n = 24/ \sqrt 144 = 2

c) P(399 < \bar x < 403 )  

= P[(399 - 400) / 2 < (\bar x - \mu \bar x) / \sigma \bar x < (403 - 400) / 2 )]

= P(-0.50 < Z < 1.50)

= P(Z < 1.50) - P(Z < -0.50)

Using z table,  

= 0.9332 - 0.3085

= 0.6247